Calculate the volume of water needed to make a 0.0150 M solution using 25.0 g of FeCl3?
1
Expert's answer
2020-06-04T10:29:05-0400

V(H2O)=m(FeCl3)/(C*Mr(FeCl3))=25/(0.015*162)=10.29 L;

Answer: 10.29L of water is needed to make  a 0.0150 M solution using 25.0 g of FeCl3